OpenStudy (anonymous):

What are the coordinates of the point 3 units from the y-axis and at a distance sqrt5 from (5, 3)?

OpenStudy (lastdaywork):

Simply use the fact - distance between (3,y) and (5,3) = sqrt(5) You'll get two such points.

OpenStudy (anonymous):

(3,y) are you looking for y ?

OpenStudy (anonymous):

let \[a(3,y) ,b(5,3)\] \[d(a,b)=\sqrt(5-3)^2+(3-y)^2=\sqrt5\]

OpenStudy (anonymous):

\[y^2-6y+8=0\]

OpenStudy (anonymous):

\[\Delta=36-4(8)=4\] \[y=2\] or \[y=4\] \[a(3,4) , a(3,2)\]